OpenStudy (anonymous):

*PLEASE CHECK MY WORK* Give the most general solutions to the equation below - PLEASE and THANKS? 2sinxcosx - sin(2x)cos(2x) = 0 Part I: Simplify the first expression using the double-angle identity for sine. Part II: Factor the left side of the equation. Part III: Solve the factored equation.

OpenStudy (anonymous):

Part I The double angle identity for sine states that sin(2x) = 2sinxcosx sin(2x) - sin(2x)cos(2x) = 0 Part II sin(2x)(1 - cos(2x)) = 0 Part III Either sin(2x) = 0 or cos(2x) = 1 For sin(2x) = 0 2x = n(pi) x = n(pi/2) For cos(2x) = 1 2x = n(pi) x = n(pi/2)

OpenStudy (anonymous):

Is this correct?

OpenStudy (unklerhaukus):

Looks good to me!

OpenStudy (anonymous):

\[\cos(2x)=1 \rightarrow 2x=0+2n \pi \rightarrow x= n \pi \] you promised to stop rushing @ilfy214 :P

OpenStudy (anonymous):

NO! You're right! Uh oh :(

OpenStudy (anonymous):

it's okay, just stop rushing and you will be fine ;)

OpenStudy (anonymous):

I will.. try! Better?\[\cos(2x)=1\]\[2x=2πn\]\[x=πn\]

OpenStudy (anonymous):

yes of course this is better, just try not simplify your steps in such a way that makes these little mistakes happen

OpenStudy (anonymous):

Okay! And thank you :D